OpenStudy (anonymous):

Please help asap! i cannot find the limit. find the limit as x approaches -infinity x / (x - (sqrt. x^2 + 7))

OpenStudy (anonymous):

i did. but i'm not getting an answer. do you know what the answer is?

OpenStudy (anonymous):

because the degree on the top and bottom is 1, then the limit is the coefficient of x, which is 1. the limit is 1.

OpenStudy (anonymous):

i think it is one half

OpenStudy (anonymous):

you are right. it was 1/2 :(

OpenStudy (anonymous):

reasoning that the denominator is going to look like \(2x\) as x goes to minus infinity, and the numerator is going to look like \(x\) and \(\frac{x}{2x}=\frac{1}{2}\)

OpenStudy (anonymous):

i don't like the sound of "was" on line class? and you put in 1?

OpenStudy (anonymous):

i thought\[\frac{x}{x-\sqrt{x^2+7}}\]

OpenStudy (anonymous):

i am assuming \[\frac{x}{x-\sqrt{x^2+7}}\]